# Prime and Co-prime Integer Matrices

### v3

@@ I Introduction
It is well-known that prime integers are the most important subjects in mathematics and have played vital roles in applications, such as cryptography, digital communications, and signal processing etc. It plays a key role in the Chinese remainder theorem (CRT) that is to reconstruct a large integer from its remainders modulo several small moduli and has many applications in cryptography, digital communications and signal processing as well [1]. For a given set of moduli, the largest integer that can be uniquely reconstructed from its remainders is smaller than the least common multiple (lcm) of all the moduli and the lcm achieves the maximal if all the moduli are pairwise co-prime, when the moduli sizes are upper bounded. This means that it is critical to have a large size set of pairwise co-prime integers.
@@
The conventional prime integers have been generalized to other prime algebraic integers in algebraic number fields with the corresponding CRT [3, 10, 11]. One example is Gaussian integers that are complex numbers with the conventional integer real and imaginary parts. Prime Gaussian integers are called Gaussian primes with simple and well-known necessary and sufficient conditions in terms of the conventional prime integers (called rational primes) [3]. In fact, an algebraic integer can be equivalently represented by an integer matrix where all the elements are the conventional integers. For a quadratic algebraic number field, see, for example, [10, 11], and in particular for a Gaussian integer of real and imaginary parts $a$ and $b$, respectively, its equivalent $2\times 2$ integer matrix representation is $\left(\begin{array}[]{cc}a&-b\\ b&a\end{array}\right)$. Thus, a Gaussian prime also means that its equivalent $2\times 2$ integer matrix (representation) is prime. Similarly, the CRT for Gaussian integers can be thought of as a CRT for $2\times 2$ integer matrices of the above special form, where Gaussian primes play the key role as well. It has recent applications in multi-channel self-reset analog-to-digital converter (SR-ADC) using modulo sampling for a complex valued bandlimited signal [18, 19].
@@
Prime integers have also been generalized to prime integer matrices in [2] where it obtains that an integer matrix is prime if and only if its determinant absolute value is a conventional prime integer. Recently, CRT has been generalized to general integer matrices of any form and any dimension in [15, 16] to reconstruct integer vectors from their vector remainders modulo several integer matrices called matrix moduli, i.e., multidimensional CRT (MD-CRT). Similar to the conventional CRT for integers, to have a large range of uniquely reconstructable integer vectors from their vector remainders, it is critical to have a large size set of pairwise co-prime integer matrices as matrix moduli. Also, co-prime integer matrices have applications in multidimensional sparse sensing [9]. Unfortunately, due to the non-commutativity of matrices, it is not obvious to construct a family of pairwise co-prime integer matrices of any dimension, although co-primality of integer matrices has been studied in [8] mostly for dimensions $2$ and $3$. Interestingly, a family of pairwise co-prime integer matrices was most recently obtained in [17] for any dimension.
@@ II Definitions and Properties of Prime Integer Matrices
$=$ $\text{lcrm}(\text{lcrm}(M_{1,1},\cdots,M_{1,L_{1}}),\cdots,$
$\quad\quad\quad\quad\text{lcrm}(M_{k,1},\cdots,M_{k,L_{k}})).$
@@
For MD-CRT [15, 16] that have applications in multidimensional signal processing, it is important to have various families of pairwise co-prime integer matrices as mentioned earlier. From Corollary 2, one can easily construct families of pairwise co-prime integer matrices by contructing matrices with different determinant absolute values that are all prime integers. With this in mind, one might want to ask what happens to the matrices with the same determinant absolute value that is a prime. In this case, any of them is prime but can they be pairwise co-prime?
@@
where $k\geq 0$ and $0\leq a_{1},...,a_{k}<\lambda_{D}$ are all integers and do not appear if $m_{0}=1$.
@@ III Connection with Gaussian Integers and Gaussian Primes
A Gaussian integer is a unit if its absolute value is $1$. It is then not hard to see that a Gaussian integer is a unit if and only if its matrix representation (or simply matrix) in (4) is a unimodular matrix. A Gaussian integer is prime if it cannot be factorized to a product two non-unit Gaussian integers. When a Gaussian integer is prime, we call it a Gaussian prime. In the following, interestingly we will see that this primality of Gaussian integers is different from that of $2\times 2$ integer matrices defined in Section II in general.
@@
We next want to see the co-primality of two matrices and two Gaussian integers. First two Gaussian integers are called co-prime if they do not have any non-unit common factor. Although the primalities of a Gaussian integer and its matrix representation are not equivalent, the co-primalities of two Gaussian integers and their matrix representations are equivalent [10].
@@
A similar integer matrix factorization can be found in [2]. Although the above connection is only studied for Gaussian integers, similar studies may apply to other algebraic integers in other algebraic number fields.

### v4

@@ I Introduction
It is well-known that prime integers are the most important subjects in mathematics and have played vital roles in applications, such as cryptography, digital communications, and signal processing etc. It plays a key role in the Chinese remainder theorem (CRT) that is to reconstruct a large integer from its remainders modulo several small moduli and has many applications in cryptography, digital communications and signal processing as well [1]. For a given set of moduli, the largest integer that can be uniquely reconstructed from its remainders is smaller than the least common multiple (lcm) of all the moduli and the lcm achieves the maximal if all the moduli are pairwise co-prime, when the moduli sizes are upper bounded. This means that it is critical to have a large set of pairwise co-prime integers.
@@
The conventional prime integers have been generalized to other prime algebraic integers in algebraic number fields with the corresponding CRT [3, 10, 11]. One example is Gaussian integers that are complex numbers with the conventional integer real and imaginary parts. Prime Gaussian integers are called Gaussian primes with simple and well-known necessary and sufficient conditions in terms of the conventional prime integers (called rational primes) [3]. In fact, an algebraic integer can be equivalently represented by an integer matrix where all the elements are the conventional integers. For a quadratic algebraic number field, see, for example, [10, 11], and in particular for a Gaussian integer of real and imaginary parts $a$ and $b$, respectively, its equivalent $2\times 2$ integer matrix representation is $\left(\begin{array}[]{cc}a&-b\\ b&a\end{array}\right)$. Thus, a Gaussian prime also means that its equivalent $2\times 2$ integer matrix (representation) is prime. Similarly, the CRT for Gaussian integers can be thought of as a CRT for $2\times 2$ integer matrices of the above special form, where Gaussian primes play the key role as well. It has recent applications in multi-channel self-reset analog-to-digital converter (SR-ADC) using modulo sampling for a complex valued bandlimited signal [18, 19, 20].
@@
Prime integers have also been generalized to prime integer matrices in [2] where it obtains that an integer matrix is prime if and only if its determinant absolute value is a conventional prime integer. Recently, CRT has been generalized to general integer matrices of any form and any dimension in [15, 16] to reconstruct integer vectors from their vector remainders modulo several integer matrices called matrix moduli, i.e., multidimensional CRT (MD-CRT). Similar to the conventional CRT for integers, to have a large range of uniquely reconstructable integer vectors from their vector remainders, it is critical to have a large set of pairwise co-prime integer matrices as matrix moduli. Also, co-prime integer matrices have applications in multidimensional sparse sensing [9]. Unfortunately, due to the non-commutativity of matrices, it is not obvious to construct a family of pairwise co-prime integer matrices of any dimension, although co-primality of integer matrices has been studied in [8] mostly for dimensions $2$ and $3$. Interestingly, a family of pairwise co-prime integer matrices was most recently obtained in [17] for any dimension. Also, more detailed examples on MD-CRT for separable and non-separable moduli can be found in [17] as well.
@@ II Definitions and Properties of Prime Integer Matrices
$=$ $\text{lcrm}(\text{lcrm}(M_{1,1},\cdots,M_{1,L_{1}}),\cdots,\text{lcrm}(M_{k,1},\cdots,M_{k,L_{k}})).$
@@
For MD-CRT [15, 16] that have applications in multidimensional signal processing, it is important to have various families of pairwise co-prime integer matrices as mentioned earlier. From Corollary 2, one can easily construct families of pairwise co-prime integer matrices by constructing matrices with different determinant absolute values that are all prime integers. With this in mind, one might want to ask what happens to the matrices with the same determinant absolute value that is a prime. In this case, any of them is prime but can they be pairwise co-prime?
@@
where $0<k<m_{0}$ and $0\leq a_{1},...,a_{k}<\lambda_{D}$ are all integers and do not appear if $m_{0}=1$.
@@ III Connection with Gaussian Integers and Gaussian Primes
A Gaussian integer is a unit if its absolute value is $1$. It is then not hard to see that a Gaussian integer is a unit if and only if its matrix representation (or simply matrix) in (4) is a unimodular matrix. A Gaussian integer is prime if it cannot be factored to a product of two non-unit Gaussian integers. When a Gaussian integer is prime, we call it a Gaussian prime. In the following, interestingly we will see that this primality of Gaussian integers is different from that of $2\times 2$ integer matrices defined in Section II in general.
@@
We next want to see the co-primality of two matrices and two Gaussian integers. First two Gaussian integers are called co-prime if they do not have any non-unit common factor. Although the primalities of a Gaussian integer and its matrix representation are not equivalent, the co-primalities of two Gaussian integers and their matrix representations are equivalent [10], which can be seen by using Bezout’s identities for both Gaussian integers and integer matrices.
@@
A similar integer matrix factorization can be found in [2]. Note that although the above connection is only studied for Gaussian integers, similar studies may apply to other algebraic integers in other algebraic number fields.